OpenStudy (anonymous):

which one is larger 500^2 or 2^500?

OpenStudy (astrophysics):

the exponent represents how many times it's being multiplied by itself

OpenStudy (astrophysics):

So for example: \[300^3 = 300 \times 300 \times 300\]

OpenStudy (astrophysics):

Now imagine \[2^{500}\] haha yikes

OpenStudy (astrophysics):

  using cm as a units for an idea